OpenStudy (anonymous):

Margie calculated that she would spend $175 on school supplies this year. She actually spent $97.50 on school supplies. What is Margie’s percent of error? 77.3% 79.5% 44.2% 4.4%

OpenStudy (anonymous):

@KamiBug @Brostep0s

OpenStudy (anonymous):

Whatcha think?

OpenStudy (anonymous):

b or c

OpenStudy (anonymous):

and you, @KamiBug ?

OpenStudy (kamibug):

Subtract 97.50 from 175, then find what percent that value is of 175. http://www.onlineconversion.com/percentcalc.htm

OpenStudy (anonymous):

175 - 97.50 = 77.5

OpenStudy (anonymous):

44.29?

OpenStudy (anonymous):

so then c?

OpenStudy (kamibug):

You got it. ;D

OpenStudy (anonymous):

that doesn't look like an available answer

OpenStudy (anonymous):

oh, precision. nevermind

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

yes because c is 44.2% but its the closest

OpenStudy (anonymous):

ok. last one :D

OpenStudy (kamibug):

Mkay

OpenStudy (anonymous):

Maxine spent 15 hours doing her homework last week. This week she spent 18 hours doing homework. She says that she spent 120% more time doing homework this week. Is she correct? Show your work to justify your decision.

OpenStudy (anonymous):

is 3 hours difference 120%

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

it is?

OpenStudy (kamibug):

3/15 *100 = 20

OpenStudy (kamibug):

So just add 20 to 100, and it's 120. :P

OpenStudy (anonymous):

so then that's the "show my work" part? 3/15 *100 = 20?

OpenStudy (anonymous):

so then can you type the steps that you did to solve the problem in order?

OpenStudy (anonymous):

that's it, d/15*100=20

OpenStudy (anonymous):

wow, that's all i have to say? "Maxine is correct because d/15*100=20

OpenStudy (kamibug):

18-15=3 3/15=0.2 0.2*100=20 20+100=120

OpenStudy (anonymous):

ok, thanks! :D

OpenStudy (anonymous):

c u later!

OpenStudy (anonymous):

medal and fan included

OpenStudy (anonymous):

bye!

OpenStudy (kamibug):

Bye :)